The initial concentration of acid HA in solution is 2.0 M. If the pH of the solution at equilibrium is 2.2, what is the percent ionization of the acid? Remember to report your answer with the correct number of significant figures Sorry, thats incorrect. Try again? [0.032]%
Answer #1

pH = -log (H+] = [+] = 10 PH pH = 2.2 (He 10- 2.2 = 6.31x10 [+]= 6.81x10²m ICE table A- lot x x HA H+ + Initial conc. (m) 2.0